OpenStudy (anonymous):

Find an equation of the parabola with vertex at (-3, 1) and focus (-1, 1)

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

\[ 4py=x^2 \]Has vertex \((0,0)\) and focus \((0,p)\)

OpenStudy (anonymous):

Yes

Directrix (directrix):

@sakigirl My ping/tag function on OS does *not work*. Nor, can I see my profile page. I asked what I thought wasTech Support to repair this two weeks ago. So far, no action. So, if you my attention, you will have to send a note.

OpenStudy (anonymous):

Translate the vertex: \[ 4p(y-1)=(x+3)^2 \]Now it has vertex \((-3, 1)\) and focus \((-3,p+1)\)

OpenStudy (anonymous):

I get (x+3)^2=4(y1)^2

OpenStudy (anonymous):

(x+3)^2=4(y-1)^2 @wio

OpenStudy (anonymous):

 Wait it is a horizontal parabola